# Data Fusion for Heterogeneous Treatment Effect Estimation with Multi-Task Gaussian Processes

### v1

@@ A.6 Confounding function
A.6.1 Variance of confounding function
@@
A.7 Evaluation metrics
The RMSE is defined as:
$\textnormal{RMSE}=\sqrt{\mathbb{E}\left[\left(\hat{\tau}({X})-\tau(X)\right)^{2}\right]}$
while the bias is defined as:
$\textnormal{Bias}=\mathbb{E}\left[\hat{\tau}({X)}-\tau(X)\right].$
where in both cases $\tau(X)$ is the true value of the CATE, while $\hat{\tau}(X)$ is the estimated value. The expectation is taken over the covariate distribution of the observational study.
A.8 Simulation Settings
A.8.1 Univariate simulation studies
We explore different simulation scenarios based on the complexity of the potential outcomes function, the confounding effect and the CATE. For each simulation setting, we compare our approach with existing ones (T-learner with GPs as regressors on the observational or experimental data, Kallus et al. [2018] 2-step method with GPs as base learners), we compute the coverage of the $90\%$ and $95\%$ credible intervals according to the values of the baseline covariate and the length of those intervals. We contrast those measures of uncertainty with the ones obtained by a T-learner approach using GPs as regressors on the observational or the experimental data. The simulations under consideration are the following:
Simulation 1: Linear Potential Outcomes, linear confounding, linear CATE
Assume that we have continuous baseline covariate $X\sim U\left[-2,2\right]$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-3-3X\right)}{1+\text{exp}\left(-3-3X\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+X+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim U\left[-2,2\right]$ and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+X+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)X,1\right)$.
Simulation 2: Nonlinear Potential Outcomes, linear confounding, nonlinear CATE
Assume that we have continuous baseline covariate $X\sim U\left[-2,2\right]$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-3-3X\right)}{1+\text{exp}\left(-3-3X\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X+X^{2}$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim U\left[-2,2\right]$ and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)X,1\right)$.
Simulation 3: Linear Potential Outcomes, linear confounding, linear CATE
Assume that we have continuous baseline covariate $X\sim U\left[-2,2\right]$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-3-3X\right)}{1+\text{exp}\left(-3-3X\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+X+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim U\left[-2,2\right]$ and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+X+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)X,1\right)$.
Simulation 4: Nonlinear Potential Outcomes, linear confounding, nonlinear CATE
Assume that we have continuous baseline covariate $X\sim U\left[-2,2\right]$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-3-3X\right)}{1+\text{exp}\left(-3-3X\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X+X^{2}$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim U\left[-2,2\right]$ and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)X,1\right)$.
Simulation 5: Linear Potential Outcomes, nonlinear confounding, linear CATE
Assume that we have continuous baseline covariate $X\sim U\left[-2,2\right]$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-3-3X\right)}{1+\text{exp}\left(-3-3X\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+X+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim U\left[-2,2\right]$ and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+X+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)\text{sin}\left(X-1\right),1\right)$.
Simulation 6: Nonlinear Potential Outcomes, linear confounding, nonlinear CATE
Assume that we have continuous baseline covariate $X\sim U\left[-2,2\right]$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-3-3X\right)}{1+\text{exp}\left(-3-3X\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X+X^{2}$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim U\left[-2,2\right]$ and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)\text{sin}\left(X-1\right),1\right)$.
Simulation 7: Linear Potential Outcomes, linear confounding, linear CATE
Assume that we have continuous baseline covariate $X\sim U\left[-2,2\right]$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-3-3X\right)}{1+\text{exp}\left(-3-3X\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+X+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim U\left[-2,2\right]$ and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+X+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)\text{sin}\left(X-1\right),1\right)$.
Simulation 8: Linear Potential Outcomes, linear confounding, nonlinear CATE
Assume that we have continuous baseline covariate $X\sim U\left[-2,2\right]$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-3-3X\right)}{1+\text{exp}\left(-3-3X\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X+X^{2}$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim U\left[-2,2\right]$ and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)\text{sin}\left(X-1\right),1\right)$.
Results: Univariate Case
According to Figure 3, the ICM approach performs equally good or outperforms all other methods. The cases where our approach shows similar performance with Kallus et al. [2018] method is when the confounding function is linear, which is expected as Kallus et al. [2018] modelling framework assumes linear functional form of confounding. In terms of coverage, both of $95\%$ and of $90\%$, we see in Figures 6 and 7 that ICM demonstrates higher coverage rates than the other two methods, indicating a more conservative approach to uncertainty. The T-learner trained on observational data shows the worst coverage rates, while in most cases the T-learner trained on experimental data achieves near optimat average coverage rate. Nevertheless, ICM is the only method demonstrating high coverage rates (even if it’s conservative) across the covariate support of the observational study, whily the T-lerner trained on observational data, in some cases (simulation 6) deteriorates outside the support of the experimental study. In terms of credible interval length, according to Figure 8 and Figure 9, the narrowest intervals are provided by the T-learner trained on observational data. In most simulation cases, the ICM offers the widest intervals, except for Simulation 1, where the T-learner trained on experimental data offers the widest. Figures 10, 11 and 12 the model fits of the two T-learner models and the Causal-ICM are demonstrated. Although there are cases where Causal-ICM has low precision in point predictions outside the support of the RCT we notice that we have appropriate variance inflation covering the true CATE function.
Figure 3: Model comparison for the univariate simulations using the distribution of RMSE obtained through 100 simulated datasets. "GP T-learner exp" and "GP T-learner obs" and stand for the T-learner approach using GPs as regressors trained on the experimental and observational data respectively and "Kallus GP" stands for Kallus et al. [2018] 2-step method with GPs as base learners.
Figure 4: Model comparison for the univariate simulations using the distribution of the Bias obtained through 100 simulated datasets. "GP T-learner exp" and "GP T-learner obs" and stand for the T-learner approach using GPs as regressors trained on the experimental and observational data respectively and "Kallus GP" stands for Kallus et al. [2018] 2-step method with GPs as base learners.
Figure 5: Model comparison for the univariate simulations using the distribution of the Variance obtained through 100 simulated datasets. "GP T-learner exp" and "GP T-learner obs" and stand for the T-learner approach using GPs as regressors trained on the experimental and observational data respectively and "Kallus GP" stands for Kallus et al. [2018] 2-step method with GPs as base learners.
Figure 6: Coverage of $95\%$ credible interval in univariate simulation settings. ICM is contrasted with T-learner approaches with GPs as regressors trained on experimental or observational data separately
Figure 7: Coverage of $90\%$ credible interval in univariate simulation settings. ICM is contrasted with T-learner approaches with GPs as regressors trained on experimental or observational data separately
Figure 8: Length of $90\%$ credible interval in univariate simulation settings. ICM is contrasted with T-learner approaches with GPs as regressors trained on experimental or observational data separately
Figure 9: Length of $95\%$ credible interval in univariate simulation settings. ICM is contrasted with T-learner approaches with GPs as regressors trained on experimental or observational data separately
Figure 10: Model fit for the T-learner trained on the observational data alone.
Figure 11: Model fit for the T-learner trained on the experimental data alone.
Figure 12: Model fit for the Causal-ICM
A.8.2 Multivariate simulation studies
@@
Simulation 2: Noninear Potential Outcomes, linear confounding, nonlinear CATE
Assume that we have five continuous baseline covariate $X_{j}\sim U\left[-2,2\right]$, where $j=1,\cdots,5$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-10-8X_{1}-8X_{2}\right)}{1+\text{exp}\left(-10-8X_{1}-8X_{2}\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+\sum_{j=1}^{5}X_{j}+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X_{1}+X_{1}^{2}+X_{2}+X_{2}^{2}$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X_{j}\sim U\left[-2,2\right]$, where $j=1,\cdots,5$, and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-\left(X_{1}+X_{2}\right)$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+\sum_{j=1}^{5}X_{j}+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)\left(X_{1}+X_{2}\right),1\right)$.
Simulation 3: Noninear Potential Outcomes, nonlinear confounding, linear CATE
Assume that we have five continuous baseline covariate $X_{j}\sim U\left[-2,2\right]$, where $j=1,\cdots,5$. The values of $X$ define the trial participation mechanism defined as $S\sim\text{Bernoulli}\left(p_{S}\right)$ where $p_{S}=\frac{\text{exp}\left(-10-8X_{1}-8X_{2}\right)}{1+\text{exp}\left(-10-8X_{1}-8X_{2}\right)}$. After selecting all trial participants we assign them randomly to treatment $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+\sum_{j=1}^{5}X_{j}+\epsilon$, for $a=0,1$, where $\tau\left(X\right)=1+X_{1}+X_{2}$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X_{j}\sim U\left[-2,2\right]$, where $j=1,\cdots,5$, and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-\left(X_{1}+X_{2}\right)$. Similarly to the trial the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+\sum_{j=1}^{5}X_{j}+U+\epsilon$, for $a=0,1$. $U$ represents the hidden confounding and we generated as $U\sim\mathcal{N}\left(\left(2A-1\right)\left(\text{sin}\left(X_{1}\right)+\text{sin}\left(X_{2}\right)\right),1\right)$.
Simulation 4: Noninear Potential Outcomes, nonlinear confounding, nonlinear CATE
@@
Results: Multivariate Case
According to Figure 13, the ICM approach outperforms all other methods. According to Table 3, ICM shows higher than optimal average coverage rates, both for $90\%$ and $95\%$ coverage, indicating a conservative view towards uncertainty. Near optimal average coverage rate is indicated by the T-leraner approach trained on experimental data, while the worst average coverage is indicated by the T-learner on the observational data. In terms of average credible interval length, according to 4, the narrowest intervals are provided by the T-learner trained on observational data. In most simulation cases, the ICM offers the widest intervals, except for Simulation 1, where the T-learner trained on experimental data offers the widest.
Figure 13: Multivariate simulation results. Boxplots of the RMSEs for each model, trained on 100 different datasets
Figure 14: Multivariate simulation results. Boxplots of the Bias for each model, trained on 100 different datasets
Figure 15: Multivariate simulation results. Boxplots of the Variance for each model, trained on 100 different datasets
Table 3: Average $90\%$ and $95\%$ Coverage rate for multivariate simulation settings
$90\%$ Coverage $95\%$ Coverage
Sim. 1 ICM ($\rho=0.9$) $99.5\%$ $100.0\%$
GP T-learner (observational) $2.5\%$ $5.5\%$
GP T-learner (experimental) $95.0\%$ $99.5\%$
Sim. 2 ICM ($\rho=0.3$) $98.5\%$ $100.0\%$
GP T-learner (observational) $60.5\%$ $77.0\%$
GP T-learner (experimental) $92.5\%$ $96.5\%$
Sim. 3 ICM ($\rho=0.3$) $99.5\%$ $100.0\%$
GP T-learner (observational) $3.0\%$ $5.5\%$
GP T-learner (experimental) $97.5\%$ $99.0\%$
Sim. 4 ICM ($\rho=0.3$) $98.5\%$ $100.0\%$
GP T-learner (observational) $69.0\%$ $78.0\%$
GP T-learner (experimental) $93.0\%$ $96.5\%$
Table 4: Average length of $90\%$ and $95\%$ credible interval for multivariate simulation settings
$90\%$ CI $95\%$ CI
Sim. 1 ICM ($\rho=0.9$) 6.05 7.20
GP T-learner (observational) 1.34 1.60
GP T-learner (experimental) 9.03 10.76
Sim. 2 ICM ($\rho=0.3$) 8.01 9.55
GP T-learner (observational) 2.31 2.75
GP T-learner (experimental) 7.64 9.10
Sim. 3 ICM ($\rho=0.3$) 10.36 12.34
GP T-learner (observational) 1.38 1.64
GP T-learner (experimental) 10.14 12.08
Sim. 4 ICM ($\rho=0.3$) 4.72 5.63
GP T-learner (observational) 1.01 1.20
GP T-learner (experimental) 4.35 5.18

### v2

@@ Appendix F Confounding function
F.1 Variance of confounding function
@@
Appendix G Multivariate simulation studies
G.1 Multivariate Data Generation Process
@@
Simulation 2: Noninear Potential Outcomes, nonlinear confounding, nonlinear CATE
@@
Figure 4: Multivariate simulation results over 100 simulated datasets. Left: Simulation setting 1: linear CATE and unobserved confounding, Right: Non-linear CATE and unobserved confounding
G.2 Multivariate Simulations: Results
Based on Figure 4, in the first simulation setting, the highest performance is achieved by the Integrative HTE and the Elastic HTE, as expected given that both models are correctly specified. Notably, the performance of Causal-ICM is comparable to these benchmarks, demonstrating its competitiveness even when the models are well-specified.
In the second simulation setting, where both the CATE and the confounding exhibit more complex, non-linear structures, the performance of models relying on correct specification deteriorates. In contrast, Causal-ICM maintains robust performance, being outperformed only by the Integrative HTE. Importantly, this superior performance of the Integrative HTE is a consequence of having a correctly specified outcome model; in realistic scenarios where model specification is uncertain or misspecified, we expect Causal-ICM to offer more reliable and consistent results.
Appendix H Extrapolation with Causal-ICM
Below we show the performance of Causal-ICM in the first univariate simulation setting, where both the CATE and the confounding functions are non-linear. The Causal-ICM can successfully produce accurate estimates of the CATE both within and outside the support of the RCT with inflated uncertainty outside the support.
Figure 5: Estimated CATE under the first univariate simulation setting with nonlinear CATE and confounding functions.
Appendix I Sensitivity to the parameter $\rho$, the sample size of the observational study, the kernel choice and the degree of overlap
Below we present a more thorough exploration of the effects of different values of $\rho$, differing observational/experimental sample sizes, the choice of kernel and the degree of overlap between the two studies for the second simulation setting of the main text where the CATE and the confounding function are non-linear.
$\rho$ RMSE
0.0 2.571
0.2 2.051
0.4 1.502
0.6 0.913
0.8 0.305
1.0 1.095
Table 2: RMSE values for different values of $\rho$, Univariate simulation 2
Kernel Mean RMSE (SD)
RBF 0.822 (0.651)
Matérn 3/2 1.315 (0.0556)
Matérn 5/2 0.968 (0.583)
Table 3: Mean RMSE for different kernels, Univariate simulation 2
Sample size Mean RMSE (SD)
$n^{(e)}=100,\;n^{(0)}=100$ 0.879 (0.484)
$n^{(e)}=100,\;n^{(0)}=500$ 0.797 (0.524)
$n^{(e)}=100,\;n^{(0)}=1000$ 0.731 (0.448)
$n^{(e)}=100,\;n^{(0)}=2000$ 0.793 (0.481)
Table 4: Mean RMSE for different sample sizes, Univariate simulation 2; Results obtained over 50 different simulated datasets.
Overlap Optimal $\rho$ Mean RMSE (SD)
Full 0.5 0.264 (0.088)
High 0.4 0.444 (0.237)
Low 0.8 0.941 (0.607)
Table 5: Mean RMSE for different degrees of overlap between the RCT and observational study for the second simualation setting of the main text (Univariate simulation 2); The degree of overlap is controlled through the coefficients of the model for study participation; Results obtained over 100 different simulated datasets.
The results above show that
  1. 1.
    $\rho$ indeed plays an important role in controlling the influence of the unobserved confounding
  2. 2.
    Causal-ICM retains strong performance for different kernels, or even as the sample size of the (confounded) observational data grows
  3. 3.
    The Causal-ICM performs well even with low overlap between the two studies, although optimal results are obtained with full overlap.
Appendix J Priors for latent processes, extentions to non-Gaussian likelihoods, and computational complexity
In the current implementation of Causal-ICM, both $u_{1}(x)$ and $u_{2}(x)$ share the same kernel hyperparameters (lengthscale, variance). However, extensions to the model do exist that allow $u_{1}(x)$ and $u_{2}(x)$ to originate from distinct Gaussian Processes, such as in the Linear Model of Coregionalization (LCM) Alvarez et al. (2012). Exploring these extensions represents a promising direction for future research.
With respect to extensions to non-Gaussian likelihoods, our theoretical results currently rely on the conjugacy of the Gaussian likelihood, which greatly simplifies the proofs. Nevertheless, we anticipate that similar results can be extended to other likelihoods that have closed-form posterior variances, such as the Student-T process. We expect these extensions to be non-trivial, however, and so leave these for future work.
For likelihoods requiring approximation methods, such as the Bernoulli likelihood, the analysis becomes significantly more intricate compared to conjugate cases. Nevertheless, we are actively exploring this direction as part of our efforts to extend causal-ICM to accommodate binary outcomes. A promising avenue involves employing a sigmoid link function to transform Gaussian Process outputs into the [0,1] interval, enabling compatibility with binary targets. While we expect comparable results in this setting, posterior approximations will be necessary. This extension represents a significant focus of our ongoing research and will play a central role in our future developments.
Last, GPs are known for their computational complexity. In our examples, we utilize the full GP as the datasets are not too large. However, their cubic computational scaling with the number of data points is indeed a well-recognized limitation when datasets are larger. To mitigate this issue, we suggest using sparse GPs as an approximation method, where we reduce computational cost by relying on a subset of the data (i.e. inducing points) to approximate the full GP model.